OpenStudy (anonymous):

Continuity of Composite Functions:Evaluate the following limit: lim sin^2(x)+6sin(x)+5/sin^2x-1 as x approaches 5pi/2

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5\pi/2}\frac{ \sin^2x+6sinx+5 }{ \sin^2x-1 }\] function rewritten.

OpenStudy (anonymous):

So far, I know that at x=1 the function is undefined so, now I need to simplify the function and try to find a removeable discontinuity that does not get me zero. This is where I'm having difficulties.

OpenStudy (dumbcow):

note: 5pi/2 is equivalent to just pi/2 to simplify things it seems by plugging in pi/2 you get ---> \[\frac{1+6+5}{1-1} = \frac{12}{0} \rightarrow \infty\]

OpenStudy (anonymous):

This is what I have so far: \[\lim_{x \rightarrow 5\pi/2}\frac{ \sin^2x+6sinx+5 }{ \sin^2x-1 }= \frac{ \frac{1-\cos2x }{2}+6sinx+5}{\frac{1-\cos2x }{2}-1}\] I got 1-cos2x/2 using the hald angle formula. Not sure if I'm on the right track or not. I'm trying to get the answer that wolfram alpha got http://www.wolframalpha.com/input/?i=factor+%28sin^2x%2B6sinx%2B5%29%2F%28sinx^2x-1%29

OpenStudy (anonymous):

factorize the numerator and denominator: \[\lim_{x->5\pi/2} \frac{ \sin^2x + 6sinx + 5 }{ (\sin^2x -1) } = \lim_{x->5\pi/2}\frac{ (sinx + 5)(sinx + 1) }{ (sinx+1)(sinx-1) }\]

OpenStudy (anonymous):

Oh I see, so the answer would be sinx+5/sinx+1=3 right?

OpenStudy (anonymous):

So if I had a function like this one\[\lim_{x \rightarrow 0^+}cot(x)\], to evaluate this one so that cot(x) doesn't become undefined as x approaches 0^+, can I do this: \[\lim_{x \rightarrow 0+}\cot(x)=\tan(x)=\sin(x)/\cos(x)=0\] So 0 would be my answer for this one, is that right as well?